OpenStudy (anonymous):

Which of the following ordered pairs lies on the graph of h(x) = -2x2? (1, -4) (1, -2) (1, 4)

OpenStudy (aravindg):

substitute the points!!

OpenStudy (anonymous):

What?

OpenStudy (aravindg):

substtitute each point in the equation

OpenStudy (anonymous):

Ok Thanks

OpenStudy (aravindg):

only one of till satisfy it ,Which is that ?

OpenStudy (anonymous):

1,4 is that right?

OpenStudy (aravindg):

for (1,4) 4=2 *(1)^2 4=2*1 4=2 ?? try again?

OpenStudy (anonymous):

(1-4)

OpenStudy (aravindg):

can you show ur working ? i will check that